OpenStudy (anonymous):

A shadow of length L is created by a(n) h =575 foot building when the sun is θ◦ above the horizon. a) Write L as a function of θ b) If θ = 20◦, find L. Answer in units of ft

OpenStudy (anonymous):

tan θ = 575 / L L = 575 / tan θ b) L = 575 / tan 20 = 1580 ft

OpenStudy (anonymous):

isnt it supposed to be multipied?